OpenStudy (anonymous):

Need help with differential equation dy/dx=(x^2-y^2)/(x^2+y^2) need to know tipe of equation bernoulli, linear, exact.

OpenStudy (jamesj):

It's none of the terms you wrote. But it is homogeneous in the sense that dy/dx = F(y/x). To solve, set u = y/x. Then y' = u + x u' and y' = (1-u^2)/(1+u^2)